Please help me solve these equations I tried to solve them my self but keep getting stuck.
solve the equations
(x-3)2 +(x+2)2=17

8 = 6
a2+a-6 a2-9



5 1 = 1
2m+6 m+1 m+3

Answer by Alan3354(69443)

Please help me solve these equations I tried to solve them my self but keep getting stuck.
solve the equations
(x-3)2 +(x+2)2=17
----------------------
If you mean:
(x-3)^2 +(x+2)^2=17
x^2-6x+9 + x^2+4x+4 = 17
2x^2 - 2x - 4 = 0
x^2 - x - 2 = 0
(x-2)*(x+1) = 0
x = 2
x = -1
----------
I don't know what those other characters mean, 8 = 6, etc.